OpenStudy (anonymous):

A building lot in a city is shaped as a 30*,-60*,-90* triangle. The side opposite the 30* angle measures 41 feet. a.) Find the length of the side of the lot opposite the 60* angle. b.)Find the length of the hypotenuse of the triangular lot. c.)Find the sine, cosine, and tagent of the 30* angle in the lot. Write your answers as decimals rounded to four decimal places.

OpenStudy (anonymous):

PLEASE HELP!!!

OpenStudy (anonymous):

 sin 30 = 41/h 1/2 = 41/h h = 2*41 (its also the answer for part b) a. using Pythagoras theorem. \[a^2 +o^2 = h^2\] we have o = 41 h = 82 now find a c. sin 30 = Opposite/ hypotenuse cos 30 = adjacent /hypotenuse tan 30 = opposite/adjacent opposite given in question hypotenuse found in part b adjacent found in part a

OpenStudy (anonymous):

can you help me further I don't understand or have a calculator lol

OpenStudy (anonymous):

sin 30 = 41/h 1/2 = 41/h h = 2*41 (its also the answer for part b) h = 82 feet a. using Pythagoras theorem. \[a^2+o^2=h^2\] we have o = 41 h = 82 \[a^2+41^2=82^2\] \[a^2+1681=6724\] \[a^2=6724 -1681\] \[a^2=5043\] \[a=\sqrt{5043}\] c. sin 30 = Opposite/ hypotenuse \[\sin 30^0 = 41/82= 1/2\] cos 30 = adjacent /hypotenuse \[\cos 30^0 =\sqrt{5043}/82 = 71.02/82\] tan 30 = opposite/adjacent \[\tan 30^0 = 41/71.02\]